# Auditing Student–AI Collaboration: A Case Study of Online Graduate CS Students

### v2

@@ 4.2. Survey Based Interviews
To examine students’ perceptions and experiences with AI-assisted coursework, we employed a survey that combined Likert-scale questions with open-ended response items. Rather than conducting separate interviews, the questionnaire itself served as the primary data collection instrument, allowing participants to provide both structured ratings and written explanations of their views.
The Likert-scale questions asked participants to evaluate their use of AI tools (e.g., frequency of use), their confidence in assessing AI-generated outputs, and their perceptions of the appropriate level of AI automation across a range of academic tasks, including writing, programming, problem solving, and studying. These items enabled quantitative analysis of general trends in students’ attitudes toward AI use in coursework.
Complementing the Likert-scale items, the survey included open-ended questions that prompted participants to elaborate on their responses. These questions asked participants to explain their reasoning behind their ratings, describe how they currently use AI tools for specific academic tasks, and articulate any concerns they have regarding accuracy, bias, over-reliance, or academic integrity when using AI systems. The open-ended responses thus functioned as survey-based interviews, providing qualitative insight into participants’ perspectives in their own words.
All responses were collected anonymously.The qualitative data from the open-ended questions were analyzed using an iterative thematicanalysis approach, beginning with open coding to identify recurring concepts,followed by axial coding to refine and organize emergent themes. This mixed-format questionnaire design allowed us to triangulate quantitative trends with qualitative explanations, strengthening the interpretability of our findings.
@@ 4.3. Ethics
ID Age Group OMSCS Specialization AI Usage Frequency
P1 30–40 Human–Computer Interaction several times per week
P2 25–30 Machine Learning Daily or almost daily
P3 25–30 Human–Computer Interaction Often
P4 18–24 Artificial Intelligence Often
P5 30–40 Human–Computer Interaction several times per week
P6 30–40 Artificial Intelligence several times per week
P7 above 40 Computing Systems several times per week
P8 30–40 Artificial Intelligence Daily or almost daily
P9 18–24 Artificial Intelligence several times per week
P10 18–24 Not specified Daily or almost daily
P11 18–24 Artificial Intelligence Daily or almost daily
P12 18–24 Not selected yet Often
P13 30–40 Not decided yet Daily or almost daily
P14 30–40 Human–Computer Interaction Daily or almost daily
P15 18–24 Undecided several times per week
P16 30–40 Computing Systems Often
P17 25–30 Artificial Intelligence Often
P18 30–40 Human–Computer Interaction several times per week
P19 25–30 Computing Systems several times per week
P20 18–24 Human–Computer Interaction several times per week
P21 30–40 Artificial Intelligence Daily or almost daily
P22 18–24 Artificial Intelligence Often
P23 30–40 Machine Learning Daily or almost daily
P24 25–30 Human–Computer Interaction Daily or almost daily
P25 25–30 Computing System Daily or almost daily
P26 30–40 Artificial Intelligence Daily or almost daily
P27 18–24 Machine Learning Daily or almost daily
P28 18–24 Artificial Intelligence / Machine Learning several times per week
P29 above 40 Artificial Intelligence Daily or almost daily
P30 18–24 Undecided several times per week
P31 25–30 Artificial Intelligence Daily or almost daily
P32 18–24 Artificial Intelligence Daily or almost daily
P33 18–24 Not specified Daily or almost daily
P34 25–30 Human–Computer Interaction several times per week
P35 18–24 Artificial Intelligence Often
P36 18–24 Artificial Intelligence Often
P37 30–40 Computing System Daily or almost daily
P38 25–30 Human–Computer Interaction Daily or almost daily
P39 25–30 Human–Computer Interaction several times per week
P40 above 40 Machine Learning Daily or almost daily
P41 18–24 Artificial Intelligence several times per week
P42 25–30 Human–Computer Interaction several times per week
P43 25–30 Human–Computer Interaction Daily or almost daily
P44 30–40 Artificial Intelligence Daily or almost daily
Table 3. Participant demographics of OMSCS students who reported at least occasional use of AI tools in coursework.
@@ 4.4. Data Analysis
Actual AI usage was operationalized using participants’ self-reported frequency of AI tool use. Responses were encoded on an ordered numerical scale, with higher values indicating more frequent AI use. The aggregate usage level was computed as the mean usage score across all participants:
$U=\frac{1}{N}\sum_{i=1}^{N}u_{i},$
where $u_{i}$ denotes participant $i$’s reported frequency of AI use.
When task-specific usage data were available, aggregate usage was computed analogously at the task level:
@@
5. User Study Result
Figure 3. Comparison of students’ desired AI automation levels and self-reported actual AI usage across twelve academic tasks (T1–T12).
Figure 4. Four-zone automation alignment map plotting desired AI automation against actual AI usage for twelve academic tasks.
@@ 5.1.1. Information Compression and Retrieval Tasks
Tasks involving summarization and information organization (T1: summarizing research articles; T11: summarizing class notes or discussions) exhibit relatively high desired automation levels alongside moderately high actual usage. These tasks are primarily extractive and supportive in nature, requiring limited original reasoning from the student. Participants appear relatively comfortable relying on AI for such tasks, likely because errors are easier to identify and correct, and inaccuracies are perceived as less consequential. Accordingly, the gap between desired automation and actual usage is comparatively small for these tasks.
@@ 5.1.2. Writing Support and Surface-Level Editing Tasks
Tasks such as revising writing for grammar and style (T4), formatting citations and bibliographies (T2), and giving feedback on draft writing (T12) show some of the highest desired automation levels across all tasks. These activities are procedural and rule-based, making them well suited for AI assistance in principle. Nevertheless, actual usage remains consistently lower than desired. This gap suggests that, despite recognizing AI’s potential utility, participants may exercise caution in practice, possibly due to concerns about correctness, formatting conventions, or adherence to academic norms.
@@ 5.1.3. Ideation and Study Support Tasks
Tasks related to brainstorming essay ideas (T3), creating personalized study plans (T6), generating flashcards or practice quizzes (T7), and recommending learning resources (T8) display some of the largest gaps between desired automation and actual usage. While participants express strong interest in AI support for these tasks, their reported usage remains comparatively low. These tasks require contextual understanding, personalization, and alignment with individual learning goals, which participants may perceive as areas where current AI tools do not yet consistently meet their expectations. The pronounced gap reflects openness to AI assistance paired with reservations about its present capabilities.
@@ 5.1.4. Communication and Socially Sensitive Tasks
Drafting professional emails to instructors or teaching assistants (T9) shows a moderate desired automation level but relatively lower reported usage compared to many other tasks. Unlike more technical activities, this gap appears to be driven less by perceptions of AI capability and more by social and interpersonal considerations. Participants may be cautious about delegating tasks that involve tone, professionalism, and accountability, preferring to retain greater human control even when AI assistance could be beneficial.
@@ 5.1.5. Analytical and High-Stakes Reasoning Tasks
Analytical tasks exhibit heterogeneous patterns. For debugging code (T5), participants report relatively high AI usage, with usage levels closely tracking desired automation. This suggests that students are already comfortable using AI tools to support code-related tasks, likely because outputs can be readily validated through testing and execution. In contrast, step-by-step quantitative problem solving (T10) shows a larger gap between desired automation and actual usage. Although participants express moderate interest in AI support for such tasks, their reported usage remains lower. These tasks demand precise reasoning and conceptual understanding, and errors may be more difficult to detect, leading participants to exercise greater caution in relying on AI assistance.
@@ 5.1.6. Four Automation Zones
Only a small number of tasks appear in regions associated with both low desired automation and low actual usage. Brainstorming essay ideas or outlining structure is the most prominent example, suggesting that participants prefer to retain human control over tasks they view as central to creativity or personal thinking.
5.2. RQ2: Primary Reasons and Concerns of using AI across academic tasks
Figure 5. Distribution of students’ reported reasons for using AI across twelve academic tasks.
Figure 6. Distribution of students’ reported concerns regarding AI use across twelve academic tasks. Percentages do not sum to 100% because the survey included an explicit “All of them” option, which could be selected by participants.
@@ 5.2.1. Writing and Revision Tasks
Writing- and revision-oriented tasks (T1: summarizing research articles; T4: revising writing for grammar and style; T9: drafting professional emails; T11: summarizing class notes or discussions; T12: giving feedback on draft writing) exhibit reason distributions that are strongly dominated by efficiency-related motivations. Across most tasks in this category, saving time accounts for the largest proportion of selected reasons, typically exceeding half of all responses. This pattern reflects students’ perception of writing-related AI use as a form of pragmatic assistance that reduces routine effort rather than fundamentally reshaping the intellectual substance of the task. In addition, improving output quality constitutes a substantial secondary share, particularly for revision-focused tasks such as T4 and T12, where AI is viewed as effective in enhancing clarity, grammatical correctness, and surface-level polish.
The combination of these motivations suggests that students conceptualize AI as a tool that complements existing writing practices rather than replaces them. Tasks in this category generally involve outputs that are readily inspectable, editable, and attributable to the student, which may further lower perceived barriers to AI adoption. As a result, students appear comfortable leveraging AI to accelerate drafting or refinement while retaining ultimate control over content and intent.
Concern distributions for writing and revision tasks are comparatively benign. For several activities (e.g., T9 and T11), None represents the largest proportion of reported concerns, indicating that many students perceive little risk in using AI for these purposes. When concerns do arise, they are most frequently associated with inaccurate information, which consistently outweighs concerns related to reduced critical thinking or plagiarism. This suggests that students’ primary vigilance lies in factual correctness rather than in broader ethical or cognitive implications. Taken together, these patterns indicate a relatively high level of trust in AI-assisted writing support, conditioned on students’ ability to verify and revise outputs as needed.
@@ 5.2.2. Ideation and Planning Tasks
Ideation and planning tasks (T3: brainstorming essay ideas or outlining structure; T6: creating a personalized study plan) demonstrate more differentiated and nuanced distributions of both reasons and concerns. For brainstorming (T3), motivations are more evenly distributed across categories, with improving output quality and reducing cognitive load together accounting for a substantial proportion of responses. This reflects students’ interest in AI as a means of generating ideas, exploring alternative structures, and overcoming initial barriers in the early stages of composition.
However, the corresponding concern distribution for T3 reveals a sharply contrasting pattern. Reduced critical thinking overwhelmingly dominates the concern profile, far exceeding all other categories. This divergence highlights a strong awareness among students of the cognitive risks associated with delegating ideation and creative reasoning to AI. While AI is valued for sparking ideas or providing structure, students appear wary of over-reliance in tasks that are closely tied to learning outcomes, originality, and intellectual ownership. The coexistence of strong perceived benefits and pronounced concerns suggests that ideation tasks occupy a particularly sensitive boundary between assistance and substitution.
In contrast, study planning (T6) exhibits a more straightforward profile. Here, saving time emerges as the dominant reason, while None constitutes the largest share of concerns. This indicates that students perceive AI-supported planning as primarily organizational rather than cognitively generative. Because study planning focuses on scheduling, resource allocation, and logistical coordination, AI assistance in this context is seen as unlikely to undermine deep understanding or critical engagement. The contrast between T3 and T6 underscores students’ ability to differentiate between tasks that involve creative reasoning and those that primarily involve coordination and efficiency.
@@ 5.2.3. Learning Support Tasks
Learning support tasks (T7: generating flashcards or practice quizzes; T8: recommending learning resources) are characterized by highly concentrated reason distributions. In both cases, saving time constitutes the clear majority of responses, with improving output quality playing a secondary role. These distributions indicate that students primarily view AI as a means of accelerating access to study materials and reducing the overhead associated with preparation and review.
Concern distributions for these tasks further reinforce this interpretation. Across both T7 and T8, None consistently represents the largest share of concerns, while accuracy- and cognition-related concerns remain comparatively minor. This suggests that students perceive these tasks as low stakes, with AI functioning as a supplementary aid rather than a substitute for core learning or evaluative processes. Because outputs such as flashcards or resource recommendations can be selectively used, ignored, or verified, students appear comfortable integrating AI with minimal hesitation. Overall, these patterns position AI as an efficiency-enhancing study aid that supports learning without threatening autonomy or understanding.
@@ 5.2.4. Technical and Quantitative Problem-Solving Tasks
Technical and analytical tasks (T5: debugging code or explaining programming errors; T10: step-by-step quantitative problem solving) exhibit the most balanced and complex distributions of reasons and concerns. For both tasks, reducing cognitive load accounts for the largest proportion of reported reasons, highlighting students’ reliance on AI to manage complexity, interpret errors, and scaffold multi-step reasoning. Improving output quality also represents a meaningful share, reflecting the importance of correctness and precision in technical contexts.
At the same time, concern distributions reveal elevated perceived risks. For both T5 and T10, inaccurate information and reduced critical thinking together constitute a substantial proportion of reported concerns, exceeding those observed in other task categories. This pattern suggests that while students value AI assistance for navigating difficult technical material, they remain acutely aware of the potential consequences of errors and the risk of superficial understanding. Unlike writing or learning support tasks, mistakes in technical or quantitative domains may be harder to detect and more costly in terms of learning outcomes. The coexistence of strong efficiency motivations and prominent concerns therefore reflects a cautious, evaluative stance toward AI use, where benefits are actively weighed against the risks of dependency and misunderstanding in high-stakes analytical contexts.
5.3. RQ3: Expected System Features and Design Principles of AI System
After participants reported their primary reasons and concerns regarding AI use across academic tasks, we administered a follow-up survey to collect open-ended responses. This resulted in 53 valid responses ($N=53$), in which participants described the features and design principles they expect AI systems to provide to address concerns about inaccurate information and hallucinations. We refer to individual respondents as R1–R53 throughout this section.
@@ 5.3. RQ3: Expected System Features and Design Principles of AI System
A dominant theme across participants’ open-ended responses concerns the need for transparency and verifiability in AI-generated outputs. Many participants emphasized that AI systems should consistently provide explicit source citations, clickable references, or other mechanisms that allow users to trace where information originates. Rather than treating AI outputs as authoritative answers, participants repeatedly framed trustworthy AI as a system that enables independent verification. As one participant explained, “Always returning a clickable link to a source so it can be easily validated” (R10). Similarly, another participant stated, “I think AI systems should always provide where it’s getting its data from so the user is able to check it on their own” (R42). These responses reflect a shared expectation that AI systems should support epistemic accountability by making their informational grounding visible, particularly in academic contexts where correctness and traceability are critical.
@@
Closely related to source transparency is the expectation that AI systems should explicitly communicate uncertainty. Participants frequently requested confidence scores, reliability metrics, or clear indicators when the system is unsure or unable to answer. Several participants suggested that AI should not attempt to mask uncertainty through fluent or confident language. For example, one participant noted, “It should have a percentage next to every response. That percentage should tell how factual AI thinks the information that it presented is” (R23), while another stated, “I think I would like if it just shares when it’s not able to answer” (R2). Importantly, some participants cautioned that confident presentation can be misleading in educational settings. As R11 argued, “Features that prioritize perceived trustworthiness undermine the nuance of AI deception,” emphasizing that productive AI use requires sustained skepticism rather than blind trust. Together, these responses suggest that transparency about uncertainty is viewed as a prerequisite for responsible and effective AI use in learning contexts.
@@
Explainability and reasoning transparency also emerged as a central design consideration. Participants expressed interest in greater visibility into how AI systems generate responses, including access to reasoning steps, intermediate logic, or the ability to intervene during response generation. One participant requested “more visibility into the inner thought process, more ability to do surgery on previous responses in a conversation” (R1), while another highlighted the importance of “clear reasoning steps, transparency about uncertainty, and real-time fact-checking” (R19). Several participants framed explainability not as a way to increase trust, but as a means of supporting critical engagement. For example, a participant described valuing systems that “show the steps of their thinking where I can intervene in the middle and reshape the direction.” These responses indicate that students prefer AI systems that function as interactive reasoning partners rather than opaque answer generators.
@@
Participants also raised strong concerns about hallucinations and misleading outputs, underscoring the need for explicit error signaling and safeguards. Many respondents requested features that flag potentially hallucinated content, provide warnings, or incorporate built-in fact-checking mechanisms. As one participant succinctly stated, “Flagging hallucinatory content. Proper warnings (not in fine print) also help” (R6). Others emphasized reducing hallucinations altogether, with one participant stating, “Reduce hallucinations” (R8), and another calling for “clear uncertainty indicators, citations linked to verifiable sources, and explanations of reasoning steps” to increase trust (R45). These responses highlight participants’ awareness of AI failure modes and their desire for systems that proactively surface limitations rather than obscuring them.
Finally, a minority of participants expressed explicit skepticism toward the use of AI in education. Rather than attempting to impose trust through persuasive or opaque design choices, participants emphasized that effective educational use of AI requires acknowledging its limitations and potential for error. As one participant succinctly reflected, “I think features that prioritize perceived trustworthiness undermine the nuance of AI deception. It seems more pragmatic to me for educational perspectives to accept that AI is a tool that typically requires some skepticism to leverage effectively.” (R11)

### v3

@@ 4.2. Survey Based Interviews
In Phase 1, participants completed a questionnaire assessing their use of AI tools and their perceptions of appropriate levels of AI automation across a range of tasks in educational contexts. Participants also reported their primary reasons and concerns regarding AI use for each task. Using the responses from Phase 1, we identified the most frequently reported concern across tasks through quantitative aggregation of participants’ ratings. In Phase 2, we conducted a follow-up survey focused specifically on this most prevalent concern. Participants were asked to describe the AI system features, design principles, or safeguards they believed would effectively address this concern in educational contexts. This second phase enabled us to move from identifying perceived risks to systematically collecting participant-informed design expectations.
@@ 4.4. Data Analysis
Actual AI usage was operationalized using participants’ self-reported frequency of AI tool use. Responses were encoded on an ordered numerical scale, with higher values indicating more frequent AI use. The aggregate usage level was computed as the mean usage score across all participants and tasks:
@@
5. User Study Results
Figure 3. Comparison of students’ desired AI automation levels and self-reported actual AI usage across twelve tasks (T1–T12).
Figure 4. Four-zone automation alignment map plotting desired AI automation against actual AI usage for twelve tasks.
@@ 5.1.1. Information Compression and Retrieval Tasks
Summarization and information organization (T1, T11) show relatively high desired automation and moderately high usage. These largely extractive, low-reasoning tasks are perceived as supportive rather than generative. Because errors are easier to detect and consequences are limited, participants are comfortable using AI, resulting in a relatively small desire–usage gap.
@@ 5.1.2. Writing Support and Surface-Level Editing Tasks
Grammar revision (T4), citation formatting (T2), and draft feedback (T12) exhibit some of the highest desired automation levels. As procedural and rule-based activities, they are well suited to AI assistance. However, actual usage remains lower than desired, suggesting caution related to correctness, formatting standards, and academic norms.
@@ 5.1.3. Ideation and Study Support Tasks
Brainstorming (T3), study planning (T6), flashcard generation (T7), and resource recommendation (T8) display some of the largest desire–usage gaps. Although participants express strong interest in AI support, usage is comparatively low. These tasks require contextualization and personalization, areas where AI is perceived as inconsistently reliable, producing openness tempered by reservations.
@@ 5.1.4. Communication and Socially Sensitive Tasks
Professional email drafting (T9) shows moderate desired automation but relatively low usage. Here, the gap appears driven less by perceived capability and more by interpersonal concerns. Participants may hesitate to delegate tasks involving tone, professionalism, and accountability.
@@ 5.1.5. Analytical and High-Stakes Reasoning Tasks
Analytical tasks vary. For code debugging (T5), usage closely aligns with desired automation, reflecting confidence in outputs that can be tested. In contrast, quantitative problem solving (T10) shows a larger gap. Despite moderate interest in AI support, participants report lower usage, likely due to the precision required and the difficulty of detecting reasoning errors.
@@ 5.1.6. Four Automation Zones
5.2. RQ2: Primary Reasons and Concerns of using AI across tasks
Figure 5. Distribution of students’ reported reasons for using AI across tasks.
Figure 6. Distribution of students’ reported concerns regarding AI use across tasks.
@@ 5.2.1. Writing and Revision Tasks
Writing- and revision-oriented tasks (T1, T4, T9, T11, T12) are strongly dominated by efficiency-related motivations. Across this category, saving time accounts for the largest share of responses, often exceeding half. Students primarily view AI as pragmatic assistance that reduces routine effort rather than reshapes intellectual substance. Improving output quality is a substantial secondary reason, particularly for revision tasks (T4, T12), where AI enhances clarity and surface-level correctness.
These patterns suggest that students see AI as complementing—not replacing—their writing. Because outputs are inspectable and editable, barriers to adoption are low, and students retain control over content and intent.
Concerns are comparatively limited. For several tasks (e.g., T9, T11), None is the most common response. When concerns arise, they focus mainly on inaccurate information, outweighing worries about plagiarism or reduced critical thinking. Overall, trust in AI writing support appears conditional on students’ ability to verify outputs.
@@ 5.2.2. Ideation and Planning Tasks
Ideation and planning tasks (T3, T6) show more differentiated patterns. For brainstorming (T3), motivations are distributed across improving output quality and reducing cognitive load, reflecting interest in idea generation and overcoming initial barriers. However, reduced critical thinking overwhelmingly dominates concerns, indicating awareness of the cognitive risks of delegating creative reasoning. Ideation thus represents a sensitive boundary between assistance and substitution.
In contrast, study planning (T6) presents a simpler profile: saving time is the dominant reason, and None is the most common concern. Students appear to view planning as organizational rather than cognitively generative, making AI use less threatening to learning.
@@ 5.2.3. Learning Support Tasks
Learning support tasks (T7, T8) exhibit highly concentrated reason distributions, with saving time clearly dominant and improving output quality secondary. Students primarily view AI as accelerating access to study materials.
Concerns are minimal: None is the largest category, and accuracy- or cognition-related worries are limited. Because outputs such as flashcards or resource recommendations can be verified or selectively used, AI is perceived as a low-stakes, supplementary aid.
@@ 5.2.4. Technical and Quantitative Problem-Solving Tasks
Technical tasks (T5, T10) display the most balanced and complex profiles. Reducing cognitive load is the leading motivation, with improving output quality also prominent, reflecting the complexity and precision required.
At the same time, concerns remain salient. Inaccurate information and reduced critical thinking together account for a substantial share, exceeding other categories—particularly inaccurate information, which is the most frequently reported concern across tasks. Students value AI for scaffolding complex reasoning but remain cautious about errors and superficial understanding. Compared to writing or general learning tasks, mistakes in technical contexts are harder to detect and potentially more consequential, leading to a more evaluative and risk-sensitive stance toward AI use.
@@
5.3. RQ3: Expected System Features and Design Principles of AI System
After participants reported their primary reasons and concerns regarding AI use across academic tasks, we administered a follow-up survey to collect open-ended responses about the design features they expected AI systems to provide to address the most prevalent concerns across tasks: inaccurate information. This resulted in 53 valid responses ($N=53$) out of the 57 participants, in which respondents described the features and design principles they expect AI systems to provide to address concerns about inaccurate information and hallucinations. We refer to individual respondents as R1–R53 throughout this section.
@@ 5.3. RQ3: Expected System Features and Design Principles of AI System
As Table 3 shows, a dominant theme across participants’ open-ended responses concerns the need for transparency and verifiability in AI-generated outputs. Many participants emphasized that AI systems should consistently provide explicit source citations, clickable references, or other mechanisms that allow users to trace where information originates. Rather than treating AI outputs as authoritative answers, participants repeatedly framed trustworthy AI as a system that enables independent verification. As one participant explained, “Always returning a clickable link to a source so it can be easily validated” (R10). Similarly, another participant stated, “I think AI systems should always provide where it’s getting its data from so the user is able to check it on their own” (R42).
@@
Closely related to source transparency is the expectation that AI systems should explicitly communicate uncertainty. Participants frequently requested confidence scores, reliability metrics, or clear indicators when the system is unsure or unable to answer. Several participants suggested that AI should not attempt to mask uncertainty through fluent or confident language. For example, one participant noted, “It should have a percentage next to every response. That percentage should tell how factual AI thinks the information that it presented is” (R23), while another stated, “I think I would like if it just shares when it’s not able to answer” (R2). Importantly, some participants cautioned that confident presentation can be misleading in educational settings. As R11 argued, “Features that prioritize perceived trustworthiness undermine the nuance of AI deception,” emphasizing that productive AI use requires sustained skepticism rather than blind trust.
@@
Explainability and reasoning transparency also emerged as a central design consideration. Participants expressed interest in greater visibility into how AI systems generate responses, including access to reasoning steps, intermediate logic, or the ability to intervene during response generation. One participant requested “more visibility into the inner thought process, more ability to do surgery on previous responses in a conversation” (R1), while another highlighted the importance of “clear reasoning steps, transparency about uncertainty, and real-time fact-checking” (R19). Several participants framed explainability not as a way to increase trust, but as a means of supporting critical engagement. For example, a participant described valuing systems that “show the steps of their thinking where I can intervene in the middle and reshape the direction.”
@@
Participants also raised strong concerns about hallucinations and misleading outputs, underscoring the need for explicit error signaling and safeguards. Many respondents requested features that flag potentially hallucinated content, provide warnings, or incorporate built-in fact-checking mechanisms. As one participant succinctly stated, “Flagging hallucinatory content. Proper warnings (not in fine print) also help” (R6). Others emphasized reducing hallucinations altogether, with one participant stating, “Reduce hallucinations” (R8), and another calling for “clear uncertainty indicators, citations linked to verifiable sources, and explanations of reasoning steps” to increase trust (R45).